A stone is thrown vertically upward it took four seconds to return to the ground how high does the stone go?


1
Expert's answer
2021-12-24T11:32:26-0500

Let's first find the time that the stone takes to reach its maximum height:


"t=\dfrac{t_{flight}}{2}=\dfrac{4\ s}{2}=2\ s."

Finally, we can find maximum height:


"h_{max}=v_0t+\dfrac{1}{2}gt^2=0+\dfrac{1}{2}\times9.8\ \dfrac{m}{s^2}\times(2\ s)^2=19.6\ m."